OpenStudy (anonymous):

A circle is shown on the coordinate plane below. What is the equation of this circle?

OpenStudy (anonymous):

What is the equation of this circle? x2 + (y + 1)2 = 4 x2 + (y – 1)2 = 2 x2 + (y + 1)2 = 2 x2 + (y – 1)2 = 4

OpenStudy (anonymous):

HELP PLLLEAAASSSE!

Parth (parthkohli):

What are the center coordinates? What is the length of the radius? Do you know the equation of a circle?

OpenStudy (anonymous):

Im completely lost on this entire section :/ could you help me ?

Parth (parthkohli):

Can you see what the center coordinates are?

OpenStudy (anonymous):

(0,1)

OpenStudy (anonymous):

right?

Parth (parthkohli):

That's right!

Parth (parthkohli):

Now can you calculate the length of the radius by counting the number of points from (0,1) to the end point?

OpenStudy (anonymous):

uhm.. 2?

Parth (parthkohli):

Yeah! Now do you have any idea of what an equation of a circle is supposed to be?

OpenStudy (anonymous):

no idea

Parth (parthkohli):

All right. Let me help you. What are two of the favourite letters of yours? Shouldn't include \(x\), \(y\) or \(r\) as they are taken.

OpenStudy (anonymous):

s and k

Parth (parthkohli):

All right then! \((x - s)^2 + (y - k)^2 = r^2\) \((s,k)\) is the center. \(r\) is the radius. x and y will just remain there.

OpenStudy (anonymous):

so whats the answer ?

Parth (parthkohli):

So, the center is (0,1) right? Our equation is something like \((x - 0)^2 + (y - 1)^2 = r^2\) Now that the radius is 2, we have something like \((x - 0)^2 + (y - 1)^2 = 2^2\) or \(x^2 + (y - 1)^2 = 4\)

OpenStudy (anonymous):

Are you 100 percent sure? & thank you for taking the time to help me! :)

Parth (parthkohli):

I am 105% sure!

OpenStudy (anonymous):

IT WAS RIGHT! :) THHANK YOOOU!